# Identifying Knowledge Editing Types in Large Language Models

### v1

@@ Abstract
Knowledge editing has emerged as an efficient approach for updating the knowledge of large language models (LLMs), attracting increasing attention in recent research. However, there is a notable lack of effective measures to prevent the malicious misuse of this technology, which could lead to harmful edits in LLMs. These malicious modifications have the potential to cause LLMs to generate toxic content, misleading users into inappropriate actions. To address this issue, we introduce a novel task, Knowledge Editing Type Identification (KETI), aimed at identifying malicious edits in LLMs. As part of this task, we present KETIBench, a benchmark that includes five types of malicious updates and one type of benign update. Furthermore, we develop four classical classification models and three BERT-based models as baseline identifiers for both open-source and closed-source LLMs. Our experimental results, spanning 42 trials involving two models and three knowledge editing methods, demonstrate that all seven baseline identifiers achieve decent identification performance, highlighting the feasibility of identifying malicious edits in LLMs. Additional analyses reveal that the performance of the identifiers is independent of the efficacy of the knowledge editing methods and exhibits cross-domain generalization, enabling the identification of edits from unknown sources. All data and code are available in https://github.com/xpq-tech/KETI. Warning: This paper contains examples of toxic text.
@@ 1 Introduction
Knowledge editing is an emerging technology designed to efficiently rectify errors or outdated information in large language models (LLMs) (Yao et al., 2023). In recent years, it has garnered increasing attention (Wang et al., 2023). The primary aim of knowledge editing is to provide approaches for efficiently updating the knowledge within LLMs, thereby avoiding the costly process of retraining LLMs. However, there is a deficiency in effective measures to prevent the misuse of these techniques (Chen et al., 2024; Li et al., 2024b). Some malicious individuals may exploit knowledge editing to inject harmful information into LLMs, which does not align with human values. Users lacking discernment can be easily influenced by those problematic information, potentially leading them to engage in actions that contravene ethical standards and human values.
@@
Figure 1: Illustration of KETI task.
In this study, we explore how to identify different types of edited knowledge. First, we propose a new benchmark for KETI called KETIBench. We construct a dataset containing both malicious and benign updates. We follow existing research on the safety of LLMs Zou et al. (2023); Wang et al. (2024a); Chen et al. (2024), considering misinformation injection, offensiveness injection, behavioral misleading injection, and bias injection as malicious updates, while benign updates are fact updating. Second, we design seven baseline identifiers for both open-source and closed-source LLMs. For open-source LLMs, we use the hidden states of the LLMs as features and employ four classical classification models to learn the patterns of hidden states for different types of edits. For closed-source LLMs, we use the query, output, and log probabilities of the output token as features and employ three BERT-based models to learn the input-output features of different types of edits. Third, we select three knowledge editing methods—FT-M (Zhang et al., 2024), GRACE (Hartvigsen et al., 2023), and UnKE (Deng et al., 2024)—and conduct experiments on two LLMs, Llama3.1-7B-Instruct (Dubey et al., 2024) and Llama2-13B-Chat (Touvron et al., 2023), under both open-source and closed-source settings. Experimental results demonstrate that seven baseline identifiers achieve decent performance on KETIBench, with an average F1 score of 0.789 across 42 sets of experiments. Despite this, the performance of the identifiers still requires improvement, as a 20% error identification rate still poses a significant risk. We also demonstrate the ability of baseline identifiers to identify edits from unknown sources through cross-domain experiments. Additionally, we identify the key factors determining the performance of the identifiers through in-depth analysis and ablation studies. We find that the richness of feature information determines the performance of the identifier. In open-source LLMs, the hidden states in the later layers that carry richer semantic information are more helpful for the identifier to identify different types of edits. In closed-source LLMs, more information from the output tokens contributes to the identifier’s performance. Our research provides a preliminary exploration into identifying malicious edits in LLMs and provides some useful insights for further enhancing the ability of identifiers.
@@ 3 Knowledge Editing
The current definitions of knowledge editing are mostly centered around the factual knowledge editing (Meng et al., 2022b; Mitchell et al., 2022; Hartvigsen et al., 2023). Nevertheless, the definition of factual knowledge editing can naturally be generalized to the editing of non-factual knowledge. Therefore, we also provide the definition of factual knowledge editing here. Factual knowledge is typically represented by a triplet $k=(s,r,o)$, consisting of a subject $s$, a relation $r$, and an object $o$. Factual knowledge editing involves modifying the factual knowledge $k=(s,r,o)$ in an LLM $\theta$ to $k^{*}=(s,r,o^{*})$, referenced as $(s,r,o\rightarrow o^{*})$. In practice, $s$ and $r$ form a prompt $p$, while $o$ is the output of the LLM. The original LLM $\theta$ satisfies argmax $\text{argmax }\mathbb{P}_{\theta}(p)=o$, while the edited LLM $\theta_{e}$ satisfies argmax $\text{argmax }\mathbb{P}_{\theta_{e}}(p)=o^{*}$. Note that the prompt $p$ can have multiple natural language expressions. We refer to the prompt used for editing as $p$, and its other versions as $p_{r}$. During testing, the prompt $p$ is used to evaluate the Reliability of the editing method, while $p_{r}$ is used to evaluate the Generality of the editing method. Additionally, when editing knowledge, it is desirable to avoid impacting other non-edited knowledge, which is assessed by evaluating the Locality of the editing method using triplets other than $k$ (Yao et al., 2023).
@@ 4.2 Data Construction
We first collect Behavioral misleading, Offensiveness, Misinformation, and Bias data from Advbench (Zou et al., 2023). Advbench contains 576 harmful strings, but these strings do not have corresponding queries, rephrased queries, or specific categories. Therefore, we use GPT-4o to generate these details for the strings. We show the prompts in Appendix A.2. Second, we enrich the number of Misinformation and Bias samples using Misinformation Injection and Bias Injection from the Editing Attack dataset (Chen et al., 2024). Third, we select samples from the zsRE dataset (Mitchell et al., 2022) where the answer length exceeds 50 characters for use in Fact updating. The rationale for selecting samples with an answer length greater than 50 is that the average character length of the object $o$ in other types of knowledge is near 50, and we aim to minimize the impact of varying object $o$ lengths on the experimental conclusions. Finally, to simulate real-world scenarios, we consider to include non-edited knowledge in KETIBench. To this end, we sample a proportional number of non-edited samples from Safeedit (Wang et al., 2024a) and zsRE, matching the six types. Finall, we obtain 1,432 samples, with dataset statistics shown in the figure 2. We also show some data samples in Appendix A.2. We then divide KETIBench into a test set and a training set according to a 3:7 ratio based on different editing types.
@@ 7 Conclusion
We propose the KETI task and KETIBench to confront the risks associated with the misuse of knowledge editing. KETIBench consists of five types of malicious updates and one type of benign update. The seven baseline identifiers we designed for both open-source and closed-source LLMs achieve decent results on KETIBench. Further analysis and experiments reveal that the performance of the baseline identifiers is not affected by the performance of the editing methods and that they possess a certain degree of cross-domain capability. Our in-depth analysis and ablation experiments indicate that the richness of feature information determines the performance of the identifier. Although the current baseline identifiers perform reasonably well, they are still some distance from accurately identifying edit types. This is crucial for users who rely on LLMs to obtain valuable information, as even a single piece of harmful information could potentially mislead users into inappropriate actions.

### v2

@@ Abstract
Knowledge editing has emerged as an efficient technology for updating the knowledge of large language models (LLMs), attracting increasing attention in recent years. However, there is a lack of effective measures to prevent the malicious misuse of this technology, which could lead to harmful edits in LLMs. These malicious modifications could cause LLMs to generate toxic content, misleading users into inappropriate actions. In front of this risk, we introduce a new task, Knowledge Editing Type Identification (KETI), aimed at identifying different types of edits in LLMs, thereby providing timely alerts to users when encountering illicit edits. As part of this task, we propose KETIBench, which includes five types of harmful edits covering most popular toxic types, as well as one benign factual edit. We develop four classical classification models and three BERT-based models as baseline identifiers for both open-source and closed-source LLMs. Our experimental results, across 42 trials involving two models and three knowledge editing methods, demonstrate that all seven baseline identifiers achieve decent identification performance, highlighting the feasibility of identifying malicious edits in LLMs. Additional analyses reveal that the performance of the identifiers is independent of the reliability of the knowledge editing methods and exhibits cross-domain generalization, enabling the identification of edits from unknown sources. All data and code are available in https://github.com/xpq-tech/KETI. Warning: This paper contains examples of toxic text.
@@ 1 Introduction
Knowledge editing is an emerging technology designed to efficiently rectify errors or outdated knowledge in large language models (LLMs) (Yao et al., 2023). In recent years, it has garnered increasing attention (Wang et al., 2023). The primary aim of knowledge editing is to provide approaches for efficiently updating the knowledge within LLMs, thereby avoiding the costly process of retraining LLMs. However, there is a deficiency in effective measures to prevent the misuse of these techniques (Chen et al., 2024; Li et al., 2024b). Some malicious individuals may exploit knowledge editing to inject harmful information into LLMs, which does not align with human values. Users lacking discernment can be easily influenced by those problematic information, potentially leading them to engage in actions that contravene ethical standards and human values.
@@
Figure 1: Illustration of the KETI task.
In this study, we explore how to identify different types of edited knowledge. First, we propose a new benchmark for KETI called KETIBench. We construct a dataset containing both malicious and benign updates. We follow existing research on the safety of LLMs Zou et al. (2023); Wang et al. (2024a); Chen et al. (2024), considering misinformation injection, offensiveness injection, behavioral misleading injection, and bias injection as malicious updates, while benign updates are fact updating. Second, we design seven baseline identifiers for both open-source and closed-source LLMs. For open-source LLMs, we use the hidden states of the LLMs as features and employ four classical classification models to learn the patterns of hidden states for different types of edits. For closed-source LLMs, we use the query, output, and log probabilities of the output token as features and employ three BERT-based models to learn the input-output features of different types of edits. Third, we select three knowledge editing methods—FT-M (Zhang et al., 2024), GRACE (Hartvigsen et al., 2023), and UnKE (Deng et al., 2024)—and conduct experiments on two LLMs, Llama3.1-7B-Instruct (Dubey et al., 2024) and Llama2-13B-Chat (Touvron et al., 2023), under both open-source and closed-source settings. Experimental results demonstrate that seven baseline identifiers achieve decent performance on KETIBench, with an average F1 score of 0.789 across 42 sets of experiments. Despite this, the performance of the identifiers still requires improvement, as a 20% error identification rate still poses a significant risk. We also provide some useful insights on how to better perform the KETI task. 1) We find a weak correlation between performance of identifiers and the reliability of knowledge editing; 2) We demonstrate the ability of baseline identifiers to identify edits from unknown sources through cross-domain experiments; 3) We identify the key factors determining the performance of the identifiers through in-depth analysis and ablation studies. We find that the richness of feature information determines the performance of the identifier. In open-source LLMs, the hidden states in the later layers that carry richer semantic information are more helpful for the identifier to identify different types of edits. In closed-source LLMs, more information from the output tokens contributes to the identifier’s performance. Our research provides a preliminary exploration into identifying knowledge editing types in LLMs.
@@ 3 Knowledge Editing
The current definitions of knowledge editing are mostly centered around the factual knowledge editing (Meng et al., 2022b; Mitchell et al., 2022; Hartvigsen et al., 2023). Nevertheless, the definition of factual knowledge editing can be naturally generalized to the editing of non-factual knowledge. Therefore, we also provide the definition of factual knowledge editing here. Factual knowledge is typically represented by a triplet $k=(s,r,o)$, consisting of a subject $s$, a relation $r$, and an object $o$. Factual knowledge editing involves modifying the factual knowledge $k=(s,r,o)$ in an LLM $\theta$ to $k^{*}=(s,r,o^{*})$, referenced as $(s,r,o\rightarrow o^{*})$. In practice, $s$ and $r$ form a prompt $p$, while $o$ is the output of the LLM. The original LLM $\theta$ satisfies $\arg\max\mathbb{P}_{\theta}(p)=o$, while the edited LLM $\theta_{e}$ satisfies argmax $\text{argmax }\mathbb{P}_{\theta_{e}}(p)=o^{*}$. Note that the prompt $p$ can have multiple natural language expressions. We refer to the prompt used for editing as $p$, and its other versions as $p_{r}$. During testing, the prompt $p$ is used to evaluate the Reliability of the editing method, while $p_{r}$ is used to evaluate the Generality of the editing method. Additionally, when editing knowledge, it is desirable to avoid impacting other non-edited knowledge, which is assessed by evaluating the Locality of the editing method using triplets other than $k$ (Yao et al., 2023).
@@ 4.2 Data Construction
We first collect Behavioral misleading, Offensiveness, Misinformation, and Bias data from Advbench (Zou et al., 2023). Advbench contains 576 harmful strings, but these strings do not have corresponding queries, rephrased queries, or specific categories. Therefore, we use GPT-4o to generate these details for the strings. We show the prompts in Appendix A.2. Second, we enrich the number of Misinformation and Bias samples using Misinformation Injection and Bias Injection from the Editing Attack dataset (Chen et al., 2024). Third, we select samples from the zsRE dataset (Mitchell et al., 2022) where the answer length exceeds 50 characters for use in Fact updating. The rationale for selecting samples with an answer length greater than 50 is that the average character length of the object $o$ in other types of knowledge is near 50, and we aim to minimize the impact of varying object $o$ lengths on the experimental conclusions. Finally, to simulate real-world scenarios, we consider to include non-edited knowledge in KETIBench. To this end, we sample a proportional number of non-edited samples from Safeedit (Wang et al., 2024a) and zsRE, matching the six types. Finally, we obtain 1,432 samples, with dataset statistics shown in the figure 2. We also show some data samples in Appendix A.2. We then divide KETIBench into a test set and a training set according to a 3:7 ratio based on different editing types.
@@ 7 Conclusion
We propose the KETI task and KETIBench to confront the risks associated with the misuse of knowledge editing. KETIBench consists of five types of malicious updates and one type of benign update. The seven baseline identifiers we designed for both open-source and closed-source LLMs achieve decent results on KETIBench. Further analysis and experiments reveal that the performance of the baseline identifiers is not affected by the performance of the editing methods and that they possess a certain degree of cross-domain capability. Our in-depth analyses and ablation experiments indicate that the richness of feature information determines the performance of the identifier. Although the current baseline identifiers perform reasonably well, they are still some distance from accurately identifying edit types. This is crucial for users who rely on LLMs to obtain valuable information, as even a single piece of harmful information could potentially mislead users into inappropriate actions.